OpenStudy (anonymous):

If the average rate of change of F on [1,3] is k, find ∫sin(t^2) dt [1,3] in terms of k

OpenStudy (jamesj):

...you mean where \[ F(t) = \int \sin(t^2) \ dt \]

OpenStudy (anonymous):

yes i think this is a trap rule Q

OpenStudy (jamesj):

or where \( F(t) = \sin(t^2) \), or what? What's your definition of F?

OpenStudy (anonymous):

the first...the integral of

OpenStudy (anonymous):

F(t)=∫sin(t^2) dt

OpenStudy (jamesj):

Well, by the Fundamental Theorem of Calculus, the rate of change of F, is the derivative dF/dt is given by \[ dF/dt = \sin(t^2) \] Now integrals find averages of things. The average of a function \( f(t) \) over an interval [a,b] is given by \[ \frac{1}{b-a} \int_a^b f(t) \ dt \] You're told that \[ \frac{1}{3-1} \int_1^3 \sin(t^2) \ dt = k \] Hence ...

OpenStudy (jamesj):

Hence what must \[ \int_1^3 \sin(t^2) \ dt \] be equal to?

OpenStudy (anonymous):

1/2 of the approximation of the integral?

OpenStudy (jamesj):

No need to approximate, none at all.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

they ask for it in terms of k

OpenStudy (jamesj):

Yes ... read again the last equations I wrote up there for you.

OpenStudy (anonymous):

1/2*k

OpenStudy (jamesj):

No

OpenStudy (anonymous):

2K

OpenStudy (jamesj):

Yes

OpenStudy (anonymous):

ok....thank you so much for walking me thru it

OpenStudy (jamesj):

ok